OpenStudy (calculusxy):

Use elimination to solve the equations: 10y + 11x = -31 -8x = 11y - 11

OpenStudy (shadowlegendx):

I abhor elimination, substitution is best!

OpenStudy (shadowlegendx):

Anyway, elimination is basically when you eliminate terms to get your answer

OpenStudy (shadowlegendx):

Thatʻs the 1st part, then the 2nd part is basically substitution

OpenStudy (calculusxy):

I dislike substitution as well but we have to learn it.

OpenStudy (calculusxy):

I had the equations: \(10y + 11x = -31\) \(-11y - 8x = -11\) I multiplied \((10y + 11x = -31)8\) \((-11y - 8x = -11)11\)

OpenStudy (calculusxy):

I then got \(8y + 88x = -248\) \(-121y - 88x = -121\)

OpenStudy (calculusxy):

Sorry I meant \(80y + 88x = -248\)

OpenStudy (shadowlegendx):

was going to say, haha

OpenStudy (shadowlegendx):

Ok, where did you go from ther

OpenStudy (shadowlegendx):

there*

OpenStudy (calculusxy):

I then had \(-41y = -369\) \(y = 9\)

OpenStudy (calculusxy):

I then plugged in 9 into where \(y\) is in the equation and solved for \(x\) \(-11y - 8x = -11\) to get \(x = -11\)

OpenStudy (shadowlegendx):

y = 9 correct

OpenStudy (calculusxy):

What about \(x\)?

OpenStudy (shadowlegendx):

That is correct a swell

OpenStudy (shadowlegendx):

As well

OpenStudy (shadowlegendx):

Good job

OpenStudy (calculusxy):

Thank you!

OpenStudy (shadowlegendx):

Looks like you have a good handle on the concept, all I had to do was check your work haha

OpenStudy (calculusxy):

I was kind of confused when the lesson was taught. So I came here to make sure whether I what was doing was correct or not.

OpenStudy (calculusxy):

*I what -> what I

OpenStudy (shadowlegendx):

Seems like you got it!